# Genome-Bench: A Scientific Reasoning Benchmark from Real-World Expert Discussions

### v1

@@ Abstract
In this short report, we present an automated pipeline tailored for the genomics domain and introduce Genome-Bench, a new benchmark constructed from over a decade of scientific forum discussions on genome engineering. Our pipeline transforms raw interactions into a reinforcement learning–friendly multiple-choice questions format, supported by 3000+ high-quality question–answer pairs spanning foundational biology, experimental troubleshooting, tool usage, and beyond. To our knowledge, this is the first end-to-end pipeline for teaching LLMs to reason from scientific discussions, with promising potential for generalization across scientific domains beyond biology. The dataset is available at https://huggingface.co/datasets/Mingyin0312/Genome-Bench.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities in domains such as mathematics liu2025understanding ; shao2024deepseekmath and programming jiang2024survey . However, despite their impressive performance on standardized reasoning benchmarks, LLMs still struggle with expert-level scientific reasoning—particularly in complex, high-stakes fields such as genomics. Unlike mathematical proofs or coding problems, which typically have clear-cut solutions and deterministic evaluation criteria, scientific reasoning often requires interpreting ambiguous observations, accounting for domain-specific nuances, and synthesizing knowledge across multiple layers of abstraction. This gap between current LLM capabilities and the level of reasoning expected from domain experts presents a major challenge for applying LLMs to real-world scientific workflows.
To bridge the gap between LLMs and human expertise, we demonstrate an end-to-end pipeline for training LLMs based on authentic scientific discussions among human experts in the field of genomics. We introduce Genome-Bench, a scientific reasoning benchmark in genomics comprising of questions and answers extracted from over a decade of archived scientific forum discussions among genome editing practitioners. These open discussions, drawn from a Google Research forum established in 2013 to explore CRISPR gene-editing technologies, reflect the nuanced, real-world problem-solving and methodological decision-making that working scientists engage in. Unlike curated benchmarks like GPQA rein2024gpqa , HLE phan2025humanity , and Lab-Bench laurent2024lab , which often emphasize standardized testing for LLMs, Genome-Bench captures authentic scientific reasoning as it unfolds in practice. It offers a unique lens into how scientists navigate common misconceptions, interpret ambiguous results, and refine experimental protocols over time. Our main contributions are summarized as follows:
@@
  • An end-to-end data pipeline for transforming expert dialogue into structured RL-compatible data. We develop an automated end-to-end pipeline that converts raw, noisy discussion email threads into structured question-answer pairs suitable for training and evaluating LLMs with reinforcement learning. This pipeline integrates LLM-assisted question-answer extraction, distractor generation, context preservation, and rigorous quality control to produce high-quality multiple-choice questions and rule-based rewards with expert-style rationales. This framework is modular and domain-agnostic, enabling scalable construction of reasoning datasets across scientific domains.
2 Related Works
@@
In biomedical and scientific domains, a variety of QA datasets have been curated to facilitate the development and evaluation of domain-specialized models. The BioASQ challenge krithara2023bioasq provides a benchmark for biomedical question answering with expert-annotated answers and linked snippets from literature tsatsaronis2015overview . In the medical domain, PubMedQA jin2019pubmedqa targets fact-based questions over biomedical abstracts, MedMCQA pal2022medmcqa spans clinical specialties, and MedQA jin2021disease and COVID-QA moller2020covidqa address disease diagnosis and pandemic-related questions, respectively. emrQA pampari2018emrqa focuses on EMR-based question answering, and MASH-QA zhu2020question explores long, multi-span answers. These datasets typically derive from formal sources such as exams, medical records, or curated biomedical corpora. More recently, LAB-Bench laurent2024lab has been introduced as a comprehensive benchmark to evaluates LLMs on real-world biology research tasks—such as literature search, protocol planning, and data analysis—offering a more practical alternative to textbook-style QA datasets. In contrast, our work introduces the first QA/MCQ dataset centered on CRISPR gene editing and constructed from authentic scientific discussions.
@@ 3 Genome-Bench
We introduce Genome-Bench dataset. An overview of the full data pipeline is shown in Figure 1.
Data Source and Significance. Genome-Bench is a new benchmark dataset designed to evaluate and advance scientific reasoning in large language models (LLMs), constructed from over a decade of CRISPR-related genome discussions discussion . The source material originates from a long-standing public mailing list established by researchers at the Broad Institute of Harvard and MIT. Spanning 2013–2023, this forum captured thousands of real-world scientific inquiries and peer responses related to gene editing. Unlike curated educational datasets or exam-derived corpora, Genome-Bench reflects how scientists actually reason in the lab: asking open-ended, ambiguous questions, proposing competing hypotheses, and resolving experimental uncertainties collaboratively. The resulting dataset comprises of high-quality question–answer pairs, touching on diverse subfields including foundational biology, experimental troubleshooting, protocol optimization, reagent selection, tool usage, and lab logistics. Because the content is grounded in expert-to-expert conversations, it captures nuanced reasoning patterns, contextual dependencies, and even common misconceptions that synthetic QA datasets often miss.
@@
Benchmark Structure. The final 3,332 questions was partitioned into a training set and a test set, comprising 2,671 and 661 questions, respectively. To enable detailed evaluation of LLM scientific reasoning, each Genome-Bench test question is annotated with two additional fields: category and difficulty. We defined seven thematic categories based on a corpus-wide analysis: Validation, Troubleshooting & Optimization, Cloning & Plasmid Construction, Gene-editing Enzyme Selection, GuideRNA Design, Screening & Library Design, Gene-editing Delivery Methods and Practical Lab Logistics. These categories reflect recurring topics in experimental CRISPR work and enable domain-specific performance breakdowns. In parallel, questions are labeled with difficulty levels—Easy, Medium, or Hard—based on linguistic structure and conceptual complexity. Heuristics used for this annotation include the presence of conditional clauses, the depth of reasoning required, and the presence of multiple plausible distractors. This stratification enables fine-grained evaluation of model robustness as cognitive load increases. We use training data for fine-tuning and test data for evaluation in the following sections.
4 Conclusion
In this paper, we present Genome-Bench, a novel benchmark for scientific reasoning in genomics, built from over a decade of authentic CRISPR forum discussions and comprising 3,300+ expertly curated multiple-choice questions. Our end-to-end pipeline transforms raw expert dialogue into structured QA data suitable for reinforcement learning, capturing the real-world complexity, ambiguity, and nuance of experimental biology—unlike synthetic or exam-based datasets. By grounding evaluation in authentic scientific discourse, Genome-Bench provides a realistic testbed for domain-specific performance. Its modular and domain-agnostic design enables extension to other scientific fields, offering a scalable framework for constructing practical, reasoning-oriented benchmarks that bring LLMs closer to integration in real-world laboratory workflows.

### v2

# Toward Scientific Reasoning in LLMs: Training from Expert Discussions via Reinforcement Learning

@@ Abstract
We investigate how to teach large language models (LLMs) to perform scientific reasoning by leveraging expert discussions as a learning signal. Focusing on the genomics domain, we develop an automated pipeline to extract trainable data and introduce Genome-Bench, a new benchmark constructed from over a decade of scientific forum discussions on genome engineering. Our pipeline transforms raw interactions into a reinforcement learning–friendly multiple-choice questions format, supported by 3000+ high-quality question–answer pairs spanning foundational biology, experimental troubleshooting, tool usage, and beyond. We fine-tune an LLM using RL with a rule-based reward signal derived from the synthetic MCQ dataset to enhance domain-specific reasoning. Our results show that reinforcement learning from scientific discussions improves model performance by over 15% compared to the base model on Genome-Bench, narrowing the gap between open-source LLMs and expert-level reasoning. To our knowledge, this is the first end-to-end pipeline for teaching LLMs to reason from scientific discussions, with promising potential for generalization across scientific domains beyond biology. Our code are available at https://github.com/mingyin0312/RL4GenomeBench.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities in domains such as mathematics liu2025understanding ; shao2024deepseekmath and programming jiang2024survey . However, despite their impressive performance on standardized reasoning tasks r1-searcher ; chakraborty2024maxmin ; guo2024embodied , LLMs still struggle with expert-level scientific reasoning—particularly in complex, high-stakes fields such as genomics. Unlike mathematical proofs or coding problems, which typically have clear-cut solutions and deterministic evaluation criteria, scientific reasoning often requires interpreting ambiguous observations, accounting for domain-specific nuances, and synthesizing knowledge across multiple layers of abstraction. This gap between current LLM capabilities and the level of reasoning expected from domain experts presents a major challenge for applying LLMs to real-world scientific workflows.
To bridge the gap between LLMs and human expertise, we demonstrate an end-to-end pipeline for training LLMs based on authentic scientific discussions among human experts in the field of genomics. As an companion effort detailed in yin2025genome , we introduce Genome-Bench, a scientific reasoning benchmark in genomics comprising of questions and answers extracted from over a decade of archived scientific forum discussions among genome editing practitioners yin2025genome . These open discussions, drawn from a Google Research forum established in 2013 to explore CRISPR gene-editing technologies, reflect the nuanced, real-world problem-solving and methodological decision-making that working scientists engage in. Unlike curated benchmarks like GPQA rein2024gpqa , HLE phan2025humanity , and Lab-Bench laurent2024lab , which often emphasize standardized testing for LLMs, Genome-Bench captures authentic scientific reasoning as it unfolds in practice. It offers a unique lens into how scientists navigate common misconceptions, interpret ambiguous results, and refine experimental protocols over time. Our findings demonstrate that reinforcement learning and mixture-of-agent training can effectively teach LLMs to emulate this kind of scientific reasoning, bridging the gap between theoretical models and the messy realities of scientific inquiry. Concretely, our main contributions are summarized as follows:
@@
  • An end-to-end data pipeline for transforming expert dialogue into RL-compatible data. We develop an automated end-to-end pipeline that converts raw, noisy discussion email threads into structured question-answer pairs suitable for training and evaluating LLMs with reinforcement learning. This pipeline integrates LLM-assisted question-answer extraction, distractor generation, context preservation, and rigorous quality control to produce high-quality multiple-choice questions. This framework is modular and domain-agnostic, enabling scalable construction of reasoning datasets across scientific domains.
  • Enhancing LLM scientific reasoning via reinforcement learning. We fine-tune the existing LLMs using a rule-based reward signal. Our approach significantly improves their performance on Genome-Bench, achieving over a 15% accuracy gain compared to instruction baselines. To further elevate the performance, we introduce a reinforcement-learned router model that dynamically selects among multiple fine-tuned expert models for each question. This “mixture-of-agents” system yields the best performance and exceeds all the state-of-the-art commercial models. Together, our methods offer a scalable and efficient approach to improving the scientific reasoning abilities of LLMs.
Paper Organization. The remainder of this paper is organized as follows. Section 2 reviews existing work on related QA benchmarks, domain-specific LLMs, and reinforcement learning for reasoning. Section 3 presents Genome-Bench and describes the data extraction pipeline. Section 4 details our RL fine-tuning methodology and introduces the expert-routing mechanism. Section 5 compares model outputs to human experts and analyzes the quality of scientific reasoning. The last section concludes the paper.
2 Related Work
@@
Bio-related QA Datasets. In biomedical and scientific domains, a variety of QA datasets have been curated to facilitate the development and evaluation of domain-specialized models. The BioASQ challenge krithara2023bioasq provides a benchmark for biomedical question answering with expert-annotated answers. In the medical domain, PubMedQA jin2019pubmedqa targets fact-based questions over biomedical abstracts, MedMCQA pal2022medmcqa spans clinical specialties, and MedQA jin2021disease and COVID-QA moller2020covidqa address disease diagnosis and pandemic-related questions, respectively. These datasets typically derive from formal sources such as exams, medical records, or curated biomedical corpora. More recently, laurent2024lab introduced LAB-Bench as a comprehensive benchmark to evaluates LLMs on real-world biology research tasks—such as literature search, protocol planning, and data analysis—offering a more practical alternative to textbook-style QA datasets. In contrast, our work introduces the first QA dataset centered on CRISPR gene editing and constructed from authentic scientific discussions.
LLMs in Scientific Domains. Large language models have increasingly been adapted for scientific reasoning by training on domain-specific corpora zhang2025scientific ; prabhakar2025omniscience ; liang2024mapping ; xie2023darwin ; pal2024domain ; cai2024sciassess ; luo2023biomedgpt ; zhang2024generalist ; chu20245 ; wang2025call . BioBERT lee2020biobert , SciBERT beltagy2019scibert , and ClinicalBERT alsentzer2019publicly are early examples of BERT-based models fine-tuned for biomedical literature and clinical notes. More recent generative models such as BioGPT luo2022biogpt , SciFive phan2021scifive , and BioBART yuan2022biobart extend this trend to sequence-to-sequence tasks. For multimodal tasks, system like GatorTron yang2022gatortron combines large-scale LLMs with structured knowledge bases. In general science domain, Galactica taylor2022galactica was trained on scientific papers and knowledge graphs to support general scientific writing and inference. In genome engineering, CRISPR-GPT huang2024crispr fine-tunes an LLM on gene-editing forum data for experimental design tasks, and BiomedGPT zhang2024generalist targets diverse multi-modal biomedical tasks.
Reinforcement Learning for LLM Reasoning. Reinforcement learning has emerged as a powerful paradigm to enhance models’ reasoning capabilities guo2025deepseekr1 ; kumar2024training ; shinn2024reflexion ; guo2025temporal ; yang2025reasonflux ; wang2024math ; r1-searcher ; zhong2024dpo ; wu2023pairwise ; wang2025call ; huang2025math . Reinforcement Learning from Human Feedback ouyang2022training utilizes preference data derived from human-annotated comparisons to optimize model behavior via policy optimization methods, notably Proximal Policy Optimization (PPO) christiano2017deep ; schulman2017proximal . To simplify training, recent work explores rule-based rewards. For example, DeepSeek-R1 devises Group Relative Policy Optimization (GRPO) shao2024deepseekmath , which forgoes the PPO value network in favor of group-wise comparisons. Building upon this, yu2025dapo introduced the Dynamic Sampling Policy Optimization (DAPO) algorithm to prevent entropy collapse and improve training efficiency. Similarly, liu2025understanding identified optimization biases in GRPO and proposed Dr. GRPO, an unbiased optimization method that improves token efficiency while maintaining reasoning performance. Recently, xiong2025minimalist combined rejection sampling and RL for reasoning. Our approach distinguishes itself from prior work by being the first to train LLMs using real scientific forum Q&A as the primary data source, rather than relying on curated corpora or manually constructed examples.
@@ 3 Genome-Bench
In this section, we introduce Genome-Bench dataset. An overview is shown in Figure 1.
Data Source and Significance. Genome-Bench is a new benchmark dataset designed to evaluate and advance scientific reasoning in large language models, constructed from over a decade of CRISPR-related genome discussions discussion . The source material originates from a long-standing public mailing list established by researchers at the Broad Institute of Harvard and MIT. Spanning 2013–2023, this forum captured thousands of real-world scientific inquiries and peer responses related to gene editing. Unlike curated educational datasets or exam-derived corpora, Genome-Bench reflects how scientists actually reason in the lab: asking open-ended, ambiguous questions, proposing competing hypotheses, and resolving experimental uncertainties collaboratively. The resulting dataset comprises of high-quality QA pairs, touching on diverse subfields including foundational biology, experimental troubleshooting, protocol optimization, reagent selection, tool usage, and lab logistics. Because the content is grounded in expert-to-expert conversations, it captures nuanced reasoning patterns, contextual dependencies, and even common misconceptions that synthetic QA datasets often miss.
@@
Benchmark Structure. The final 3,332 questions was partitioned into a training set and a test set, comprising 2,671 (80%) and 661 (20%) questions, respectively. To enable detailed evaluation of LLM scientific reasoning, each Genome-Bench test question is annotated with two additional fields: category and difficulty. We defined seven thematic categories based on a corpus-wide analysis: Validation, Troubleshooting & Optimization, Cloning & Plasmid Construction, Gene-editing Enzyme Selection, GuideRNA Design, Screening & Library Design, Gene-editing Delivery Methods and Practical Lab Logistics. These categories reflect recurring topics in experimental CRISPR work and enable domain-specific performance breakdowns. In parallel, questions are labeled with difficulty levels—Easy, Medium, or Hard—based on linguistic structure and conceptual complexity. Heuristics used for this annotation include the presence of conditional clauses, the depth of reasoning required, and the presence of multiple plausible distractors. This stratification enables fine-grained evaluation of model robustness as cognitive load increases. We use training data for fine-tuning and test data for evaluation in the following sections.
4 Enhancing Scientific Reasoning via RL and Expert Routing
Figure 2: Benchmarking LLMs’ biological reasoning capabilities on Genome-Bench. (a) Overall accuracy across Genome-Bench is presented for 21 different LLMs. Smaller models, such as Qwen2.5-7B, show substantial performance improvements when fine-tuned with reinforcement learning (RL), achieving results comparable to leading models like DeepSeek-R1. The RL-Router model further enhances accuracy by dynamically assigning each question to one of four specialized fine-tuned models, selecting the one most likely to provide a correct answer. (b) Comparative accuracy of models across seven biological categories reveals variability in performance. (c) Accuracy is detailed by question difficulty levels—easy, medium, and hard.
4.1 Reinforcement Fine-Tuning for Scientific Reasoning
Algorithm and Experimental Setup. To instill stronger reasoning abilities in LLMs, we fine-tune different open-source models on Genome-Bench using reinforcement learning (RL) and benchmark their performance alongside leading commercial models. The models we evaluate fall into two categories: Base models, which refer to open-source instruction-tuned models, and Commercial models, which represent the current state-of-the-art. Details of these models are summarized in Table 1. We refer to the fine-tuned versions of the Base models as RL models.
Specifically, we instantiate RL with Group Relative Policy Optimization (GRPO) shao2024deepseekmath , a sample-efficient alternative to Proximal Policy Optimization (PPO) schulman2017proximal that replaces the traditional value network with group-wise comparisons (check Appendix C). In GRPO, for each input question $q$, the model generates a group of $G=4$ candidate outputs from the current policy. A reward is assigned to each output based on its correctness and formatting, as detailed below.
Reward Function Design. We employ rule-based reward design lambert2024t ; liu2025understanding to align with two objectives of scientific question answering: producing the correct answer and providing extractable format. If the model’s response contains a clearly delimited <explanation> followed by a <answer> section, it earns +1 format reward. If the model selects the correct multiple-choice answer (identified by the <answer> tag), it receives +2 correctness reward. Otherwise, it receives 0. A well-structured but incorrect answer receives some credit, and a correct but malformed response avoids full reward. Importantly, the enforced <explanation>+<answer> format encourages model to ground its answer selection based on its reasoning.
4.2 Experiment Results
Figure 2 presents a comprehensive evaluation of 21 LLMs on the Genome-Bench test data, revealing key insights into how reinforcement learning (RL) improves scientific reasoning in genomics. In Figure 2(a), we observe that small open-source models—such as Qwen2.5-7B and Llama3.1-8B—exhibit substantial performance gains (>15%) after RL fine-tuning, in some cases surpassing commercial baselines like Claude-3.5 and Gemini-2.0. This improvement is also consistently observed in smaller model Qwen2.5-1.5B. This confirms the effectiveness of reinforcement learning in boosting reasoning capabilities in science domains, especially for models with limited pretraining in scientific contexts. Notably, this finding stands in contrast to prior RL applications in math and coding, which typically involve problems with deterministic answers.
We further devise the RL-Router model (details in Section 4.3) that achieves the highest accuracy overall (81.07%), surpassing all individual RL models and even outperforming state-of-the-art commercial systems like DeepSeek-R1. This highlights the advantage of dynamically selecting from a pool of expert models rather than relying on a monolithic architecture. In Figure 2(b), accuracy across seven biological categories reveals that the RL-Router maintains strong performance uniformly across domains—ranging from experimental troubleshooting to gene-editing enzyme selection—whereas individual models show domain-specific strengths and weaknesses. Finally, Figure 2(c) shows that while all models exhibit a decline in accuracy as question difficulty increases, the RL-Router demonstrates strong robustness across easy, medium, and hard tiers. The complete evaluation is deferred to Appendix D. Together, these results suggest that RL and expert routing can significantly narrow the gap between open-source LLMs and expert-level scientific reasoning.
Ablation Analysis: Effectiveness of RL Fine-Tuning. Figure 3(a) shows comparisons across base, SFT, and RL-tuned versions of five LLMs. Reinforcement learning yields significant gains over SFT for all models (>10%). Besides, RL keeps consistent performance for both train and test, while SFT exhibits overfitting for small (1.5B) model. Figure 3(b) further demonstrates PassK evaluation, where the 7B and 1.5B RL models reveal different patterns as K increases. For 7B model, RL method consistently outperforms base models as K increases, while for the 1.5B model, the performances between RL and base models cross as K increases. This finding suggest that the effectiveness of reinforcement learning in helping models acquire and generalize new knowledge might be model-size dependent. Larger models, such as the 7B variant, appear capable of leveraging RL to refine both accuracy and diversity, while smaller models may lack sufficient capacity to do so. This highlights an important design consideration when applying RL to models of varying scales.
Is reinforcement learning enabling accurate reasoning? To assess whether RL improves not only answer accuracy but also the quality of reasoning, we analyzed 10 questions (Appendix H) for which the model initially produced incorrect answers but answered correctly after RL fine-tuning. Our analysis reveals a nuanced picture of RL’s impact on reasoning capabilities. On one hand, we observe promising improvements in reasoning paths, such as in question 23, where the RL-tuned model demonstrates correct identification of potential issues raised by the user and provides well-structured logical analysis. However, the results also highlight persistent challenges. For instance, question 7 shows that while RL helps the model arrive at the correct answer, the underlying reasoning remains flawed or incomplete. Similarly, question 27 presents cases where RL successfully guides the model toward the correct answer, but the reasoning remains vague, often simply asserting that a particular choice is best based on context without providing substantive justification. These findings suggest that while RL can effectively improve answer accuracy, ensuring that the model consistently develops sound reasoning processes remains an open challenge and requires further delicate investigation.
Figure 3: Ablation Analysis Comparing SFT vs RL. (a) Comparison of model accuracy across different training stages—Base, SFT (Supervised Fine-Tuning), and RL on both training and test splits. Notably, RL improves generalization on the test set, which is consistent with the performance on the training data. Besides that, RL significantly outperforms SFT. This highlights the advantage of reinforcement learning in aligning model behavior with expert reasoning signals. Moreover, the small model Qwen2.5-1.5B exhibits signs of overfitting during SFT. For instance, the training accuracy under SFT reaches 68.84%, but the test accuracy drops to just 50.23%. RL does not exhibit the similar issue. (b) Pass@K evaluation on 7B and 1.5B models. For 7B model, RL method consistently outperforms base models as K increases, suggesting that it acquires new knowledge, as evidenced by the persistent performance gap. In contrast, for the 1.5B model, the performance gap narrows as K increases. This indicates that RL primarily improves the robustness of the output distribution rather than introducing fundamentally new capabilities—likely due to the limited capacity (1.5B).
4.3 Router Model with Reinforcement learning
RL-Router: Dynamic Model Selection via Reinforcement Learning. While reinforcement learning enhances individual models, no single model excels at all question types. As shown in Figure 4(b), the pairwise matrix reveals substantial specialization across the four RL-tuned experts. For instance, among the total 661 test problems, the Llama-3.1-8B-RL model correctly answered 87 questions that Gemma2-9B-RL failed, while Qwen2.5-7B-RL correctly answered 96 questions missed by Llama-3.1-8B-RL. These discrepancies underscore the diversity in model strengths and suggest strong complementarity among the experts—indicating untapped potential if they can be coordinated.
To exploit specialization, we propose a novel “mixture-of-agents” system: an RL-trained router that dynamically selects the most suitable expert model to answer each question. The router, implemented using Qwen2.5-7B-Instruct, receives the question as input and predicts an index in 1, 2, 3, 4 corresponding to one of the experts as shown in Figure 4(a). The router is trained via GRPO using a simple reward signal: it receives +1 if the selected expert produces the correct answer, and -1 otherwise. Pre-generated responses from the experts are used to accelerate training. Over time, the router learns to recognize patterns in question content and domain that correlate with each expert’s strengths.
Experiment Results. Figure 4(c) shows that the RL-Router consistently outperforms all individual RL-tuned models, achieving a new state-of-the-art accuracy of 81.1% on the Genome-Bench test set. Such an improvement arises from the router’s ability to dynamically allocate questions to the expert model best suited for each query—as further verified in the center panel of Figure 4(c), where routed questions consistently yield higher accuracy than non-routed ones. The routing distribution is provided in Figure 13. Together with Figure 2(a), the RL-Router also exceeds the performance of top-tier commercial systems such as DeepSeek-R1. Beyond aggregate accuracy, Figures 2(b) and 2(c) demonstrate that routing leads to robust improvements across all the scientific categories and difficulty levels. This adaptability underscores the router’s ability to make fine-grained, context-aware decisions comparing to the existing models.
Figure 4: Reinforcement Learning for Model Routing. (a) Overview of the routing workflow: for each Genome-Bench question, the RL-trained router selects one of four fine-tuned expert models to generate a response, receiving a positive or negative reward based on correctness. (b) Pairwise matrix among the RL-tuned experts: diagonal entry $(i,i)$ counts the number of questions only model $i$ answered correctly; off-diagonal entry $(i,j)$ counts questions where model $i$ was correct and model $j$ was not. (c) Comparative performance of the RL router. Left: overall Genome-Bench accuracy versus its underlying expert models; Center: accuracy breakdown on routed vs. non-routed questions; Right: ablation results showing router accuracy when selecting among base (non-RL) models, demonstrating gains from routing even without expert fine-tuning.
Ablation Analysis: Understanding Routing Efficacy. We conduct ablations to isolate the router’s impact. First, we evaluate a router trained over base (non-RL) models (Figure 4(c),left;Figure 14). Even without RL-enhanced experts, the router improves overall accuracy, highlighting the inherent benefit of expert selection. Second, we experiment with a smaller router (Qwen2.5-1.5B, Figure 12). This underpowered router converges to favoring a single expert, failing to learn nuanced routing policies. As a result, accuracy gains are limited. This phenomenon suggests that routing itself is a learning task requiring sufficient model capacity and that deploying a capable LLM as the router is crucial for harnessing the full benefits of specialization. More details are deferred to Appendix F.
In summary, our combination of reinforcement fine-tuning and RL routing yields state-of-the-art performance on Genome-Bench. This framework offers a scalable path toward scientific LLMs that both reason accurately and know when to defer to the most competent expert.
5 Human Evaluation Analysis
Figure 5: Human Evaluation and Analysis of Model Reasoning. Accuracy comparison on 35 expert-curated questions, evaluated across Qwen2.5-7B-RL, DeepSeek-R1, the RL-Router model, and human experts. The RL-Router model achieves the highest model accuracy (67.6%), approaching the performance of human experts (73.3%). Qualitative analysis of RL-Router model-generated explanations compared to human-provided reasoning. Each example is categorized as Correct Reasoning, Partially Wrong Reasoning, or Completely Wrong Reasoning based on the scientific soundness and alignment with the ground-truth rationale.
Human Evaluation Design. To rigorously assess the LLM’s scientific reasoning ability, we conducted a controlled human expert evaluation using a curated subset of questions from Genome-Bench. We selected 35 challenging and diverse items spanning a range of biological topics with difficulty levels being Medium or Hard. These questions were chosen to reflect realistic problems encountered in genome engineering practice. We recruited three domain experts in CRISPR and molecular biology. Each expert was instructed to complete the 35 questions in 60 minutes without the use of external sources. The format contains a justification enclosed in <explanation> tags, followed by a selected option in <answer> tags. This design not only facilitated direct comparison between human and models, but also preserved interpretable reasoning traces. The resulting human annotations serve as a human expert reference set, enabling both quantitative benchmarking and qualitative assessments.
We further posed the same 35-question set to three different models: 1. Qwen2.5-7B-RL, our strongest single model fine-tuned using RL; 2. DeepSeek-R1, a top-performing baseline commercial LLM; 3. The RL-Router model, our mixture-of-agents system. Each model was queried in three independent runs, and all model outputs adhered to the <explanation>/<answer> structure.
Quantitative Results. In Figure 5(a), the average human expert accuracy across the 35 questions was 73.3%, establishing a baseline for expert-level reasoning. Among the models, the RL-Router achieved 67.6% accuracy, the highest among all tested LLMs and the closest to human expert performance with the gap 5.7%. As a comparison, DeepSeek-R1 scored 65.7%, and the Qwen2.5-7B-RL achieved 61.9%. These results demonstrate that our reinforcement-tuned routing approach significantly narrows the performance gap between open-source LLMs and human experts, highlighting the effectiveness of RL training using structured discussion data in combination with routing strategies.
Qualitative Analysis. Beyond final answer accuracy, we conducted a fine-grained qualitative evaluation of the model’s reasoning chains in Figure 5(b)-(d). Each model-generated explanation was evaluated by the human expert and categorized into one of three reasoning types.
Correct Reasoning 5(b): The model selected the correct answer and provided a scientifically coherent explanation that mirrored expert reasoning. For instance, in a question regarding the use of GFP as a transfection control, the model and human both identified co-transfection with a GFP vector as the correct strategy and provided justifications rooted in standard laboratory practice. These cases illustrate that the model has internalized domain-appropriate problem-solving logic.
Partially Wrong Reasoning 5(c): The model reached the correct answer, but its explanation contained factual inaccuracies, omissions, or speculative logic. For example, in a question about sgRNA screening outcomes, the model selected the right answer (“off-target effects”) but mischaracterized the biological mechanism. These instances suggest that the model may arrive at the correct solution by chance or by leveraging surface-level cues, rather than engaging in deep causal reasoning.
Completely Wrong Reasoning 5(d): The model’s explanation failed to reflect valid scientific reasoning, even when the final answer happened to be correct. As an example, when asked about the impact of guide RNA self-targeting on homologous recombination, the model claimed the process would be completely blocked—contradicting to its own selection of the answer “e”, which correctly noted only a reduction in recombination frequency. These failures expose the fragility of LLM reasoning fidelity and the limitations of current training in capturing causal understanding.
Discussion and Implications. This evaluation yields two core findings. First, reinforcement-trained LLMs are rapidly approaching expert-level accuracy on complex, domain-specific scientific questions. The RL-Router model in particular demonstrates that careful integration of domain-specific data, structured reasoning supervision, and dynamic expert selection can bring open-source LLMs within striking distance of human expert performance. Second, and equally important, reasoning fidelity remains an open challenge. While models increasingly get the right answers, their reasoning pathways may still diverge significantly from expert logic. This presents risks for downstream scientific applications, where valid justifications are as critical as correct predictions. Our findings reinforce the need to supervise the reasoning process explicitly, not just the answer outcome. Collectively, this evaluation highlights the significant progress made by RL-trained LLMs in scientific reasoning. Figure 5 serves both as a testament to current capabilities and as a diagnostic lens, revealing areas where models still fall short in the step-by-step reasoning that characterizes human scientific thought.
6 Conclusion and Limitations
We present an end-to-end framework for LLMs to perform scientific reasoning by training on real-world expert discussions. We introduce Genome-Bench, a benchmark comprising over 3,000 multiple-choice questions from CRISPR-related forums, and apply reinforcement fine-tuning with a structured reward signal to achieve over 15% accuracy gains compared to existing models. Our RL-Router model, which dynamically selects among specialized experts, further attained 81.1% accuracy on Genome-Bench. While RL improves answer accuracy, qualitative analysis reveals that reasoning remains imperfect, with some explanations lacking causal or scientific rigor. Future work may explore hybrid reward mechanisms that penalize inconsistent reasoning, leverage human-in-the-loop critiques, or align LLM explanations with mechanistic scientific principles. We leave these as future works.